### v1

@@ 1 Introduction
Calculus courses serve as gateways to advanced study in virtually all STEM disciplines. In these courses, students are expected not only to obtain correct numerical or algebraic results but also to formulate models, select appropriate techniques, justify steps, and communicate reasoning with symbols, diagrams, and words. Research on learning and assessment consistently shows that such open-ended work — showing the path, not just the endpoint — better reflects the knowledge and practices we intend to cultivate, supports transfer, and makes students’ conceptions visible for feedback and instruction [1, 2, 3]. Yet in large-enrollment settings, logistical pressure often pushes assessment toward closed-answer formats that can be machine-graded at scale, narrowing what is measured and learned.
@@
Closed-answer technologies have a long history, from mechanical multiple-choice devices to modern web-based systems that grade numeric responses with tolerances, check algebraic equivalence, and match short strings [4, 5, 6, 7]. These approaches can deliver high scoring consistency when the space of correct answers is well defined. However, they struggle to capture the multi-step reasoning, representation shifts, and communication quality that matter in mathematics learning, especially in calculus where procedures (e.g., differentiation techniques) intertwine with concepts (e.g., limits, continuity, and the meaning of the derivative) and with modeling and interpretation [8, 9]. Automatic Short Answer Grading (ASAG) occupies a middle ground by recognizing paraphrase and semantic equivalence in brief textual responses, typically with machine-learning models [10, 11, 12]. Yet ASAG generally does not address the mixed-media nature of authentic calculus work: symbolic derivations, structured chains of reasoning, and sketches of graphs or geometric configurations.
@@
Recent advances in large language models (LLMs) and multimodal systems have re-opened the possibility of grading assistance for open-ended work at scale [13, 14, 15]. Across various fields of education, LLMs can evaluate and generate academic work and support research workflows [16, 17, 18, 19]; in STEM education, researchers have begun to explore their use for solving, generating, and assessing problems [20, 21, 22, 23]. Our own prior studies suggest both promise and limits. First, benchmarking GPT-4 on ASAG showed performance comparable to earlier hand-engineered systems, with the practical advantage of no task-specific training and, in some cases, grading without reference solutions [24]. Second, in handwritten undergraduate mathematics, we proposed a framework to estimate the reliability of short-answer grades and found that recognition of mathematical notation is a key bottleneck; scanning and transcribing whole pages before extracting answers can mitigate errors [25]. Third, in open-ended physics and chemistry exams, end-to-end pipelines revealed that layout and recognition quality strongly affect downstream grading, and that human-in-the-loop routing remains essential due to occasional high-confidence errors [26, 27, 28]. These studies also demonstrated the value of psychometric instrumentation, using Item Response Theory (IRT) and related tools to quantify when and how AI grades can be trusted [29, 30, 31, 32, 33].
@@
A central technical and methodological challenge for mathematics education is that student work is genuinely mixed-media. Mathematical handwriting must be transcribed with structure preserved; diagrams and graphs need faithful descriptions; and printed headers and rubric anchors must remain aligned through scanning and registration. Classical OCR is strained by mathematics and layout variability [34, 35, 36]; while specialized tools can help for formulas [37], multimodal LLMs promise more integrated “see-and-grade” pathways [15]. At the same time, educational use requires calibrated uncertainty: instructors need to know when automation is reliable without first grading everything by hand. Human-in-the-loop designs, i.e., routing clear, routine cases to automation and flagging uncertain or pedagogically rich cases for expert review, offer a pragmatic compromise [38, 39]. To be trustworthy, such systems must report well-calibrated confidence, meet explicit quality targets, and support appeals and transparency obligations that many jurisdictions now expect for high-stakes educational AI [40, 41].
@@
Our design draws on prior evidence and seeks mathematics-specific insight. Building on our ASAG benchmark [24] and reliability framework for handwritten mathematics [25], we treat confidence calibration and routing as first-class objects: we integrate model-based signals (e.g., log probabilities, self-consistency), rubric alignment, and transcript/recognition quality into a calibrated probability of correctness, then set operating points that respect course and departmental constraints [33]. In parallel, we consider practical layout affordances linked to recognition and rubric alignment, echoing findings from physics and chemistry that small design choices can have outsized downstream effects [27, 28]. Throughout, we frame validity in terms of what matters for calculus learning: visibility into reasoning, representation changes (algebraic steps, graphs), and communication quality, not only terminal answers [9, 8].
@@ 2.2 Sample
The student work was graded by teaching assistants (TAs) based on four to six rubric items per problem. For our study, we scanned 349 exams; Figure 1 shows a short excerpt as an example of graded student work. We were able to remove the red, pink, and green grading marks using the image-editing package OpenCV [42], as students were not allowed to use these colors, and we combined the pages for each problem, or, in case of problem 5, problem part; Figure 2 shows an example.
@@ 2.3 AI Grading
We accessed OpenAI models via Azure AI Services [43] under the university’s privacy-preserving contractual framework. Specifically, we used the multimodal reasoning model GPT-5 [44] (model version 2025-08-07) hosted in a Swedish data center. The code, including the prompt, is available at https://gitlab.ethz.ch/ethel/mathexam/-/blob/main/grade.py.
@@ 2.4 Confidence Filters
Generative AI will always produce an answer, regardless of whether that answer is reliable or not.. Thus, beyond computing AI-based scores, an equally important task is quantifying how much we should trust each score [33]. “Confidence” here differs from conventional quality metrics: in production use the ground truth is unknown, so confidence must be inferred from information available at decision time (the AI’s own outputs and model-based expectations). To do so, we draw on classical and Bayesian statistics.
@@ 2.4.1 Partial-credit Threshold
Evidence from earlier studies indicates that AI graders tend to be conservative — more prone to mark correct work as incorrect (false negatives) than the reverse (false positives) [33]. A simple, conservative safeguard is therefore to binarize partial credit using a threshold $t$ for “correctness.” For filtering purposes, any student-item with AI score below $t$ (i.e., incorrect or “insufficiently correct”) is flagged for human review, while only items at or above $t$ are considered for auto-acceptance (subject to the risk screen below).
@@ 2.4.2 Risk Threshold
where $\theta_{i}$ denotes the latent ability of student $i$, and $a_{j}$ and $b_{j}$ are the discrimination and difficulty of part $j$, respectively [45, 46].
Let $s_{ij}\in[0,1]$ be the AI’s normalized score for the same student-item. We define the risk of accepting that AI judgment as the absolute deviation between observed and expected [47]:
@@ 4 Discussion
The results are moderately promising but not perfect. Production-ready results can only be achieved for about 30% of the grading load, which is less than in our earlier studies of physics [33] and chemistry exams [28]. This might be improved by some practical, incremental adjustments that do not alter the mathematical substance of the exam:
@@ Acknowledgements
Ethics Approval
The study was governed by ETH Ethics Board (IRB) protocol 2023-N-286.
@@
Code Available
The code is available open-source at https://gitlab.ethz.ch/ethel/mathexam.

### v2

@@ 1 Introduction
Calculus courses serve as gateways to advanced study in virtually all STEM disciplines. In these courses, students are expected not only to obtain correct numerical or algebraic results but also to formulate models, select appropriate techniques, justify steps, and communicate reasoning with symbols, diagrams, and words. Research on learning and assessment consistently shows that such open-ended work — showing the path, not just the endpoint — better reflects the knowledge and practices we intend to cultivate, supports transfer, and makes students’ conceptions visible for feedback and instruction [bloom56, laverty12b, offerdahl2019formative]. Yet in large-enrollment settings, logistical pressure often pushes assessment toward closed-answer formats that can be machine-graded at scale, narrowing what is measured and learned.
@@
Closed-answer technologies have a long history, from mechanical multiple-choice devices to modern web-based systems that grade numeric responses with tolerances, check algebraic equivalence, and match short strings [petrina2004sidney, suppes1966arithmetic, sangwin2007assessing, kortemeyer08]. These approaches can deliver high scoring consistency when the space of correct answers is well defined. However, they struggle to capture the multi-step reasoning, representation shifts, and communication quality that matter in mathematics learning, especially in calculus where procedures (e.g., differentiation techniques) intertwine with concepts (e.g., limits, continuity, and the meaning of the derivative) and with modeling and interpretation [teodorescu2013new, meltzer2005relation]. Automatic Short Answer Grading (ASAG) occupies a middle ground by recognizing paraphrase and semantic equivalence in brief textual responses, typically with machine-learning models [leacock2003c, zhang2022automatic, ahmed2022deep]. Yet ASAG generally does not address the mixed-media nature of authentic calculus work: symbolic derivations, structured chains of reasoning, and sketches of graphs or geometric configurations.
@@
Recent advances in large language models (LLMs) and multimodal systems have re-opened the possibility of grading assistance for open-ended work at scale [chatgpt, gpt4, gpt4o]. Across various fields of education, LLMs can evaluate and generate academic work and support research workflows [meyer2023chatgpt, kasneci2023chatgpt, tschisgale2023integrating, kieser2023educational, tolstykh2024beyond, campbell2025applying]; in STEM education, researchers have begun to explore their use for solving, generating, and assessing problems [yeadon2024impact, sperling2024artificial, polverini2024understanding, kortemeyer23ai]. Our own prior studies suggest both promise and limits. First, benchmarking GPT-4 on ASAG showed performance comparable to earlier hand-engineered systems, with the practical advantage of no task-specific training and, in some cases, grading without reference solutions [kortemeyer2024performance]. Second, in handwritten undergraduate mathematics, we proposed a framework to estimate the reliability of short-answer grades and found that recognition of mathematical notation is a key bottleneck; scanning and transcribing whole pages before extracting answers can mitigate errors [liu2024ai]. Third, in open-ended physics and chemistry exams, end-to-end pipelines revealed that layout and recognition quality strongly affect downstream grading, and that human-in-the-loop routing remains essential due to occasional high-confidence errors [kortemeyer24aigrading, kortemeyer2024grading, cvengros2025assisting]. These studies also demonstrated the value of psychometric instrumentation, using Item Response Theory (IRT) and related tools to quantify when and how AI grades can be trusted [lord2008statistical, pawl2013, scott2006, kortemeyer2016psy, kortemeyer2025assessing].
@@
A central technical and methodological challenge for mathematics education is that student work is genuinely mixed-media. Mathematical handwriting must be transcribed with structure preserved; diagrams and graphs need faithful descriptions; and printed headers and rubric anchors must remain aligned through scanning and registration. Classical OCR is strained by mathematics and layout variability [mori1992historical, okamura1999handwriting, wang2021]; while specialized tools can help for formulas [mathpix], multimodal LLMs promise more integrated “see-and-grade” pathways [gpt4o]. At the same time, educational use requires calibrated uncertainty: instructors need to know when automation is reliable without first grading everything by hand. Human-in-the-loop designs, i.e., routing clear, routine cases to automation and flagging uncertain or pedagogically rich cases for expert review, offer a pragmatic compromise [de2020case, alsalmani23a]. To be trustworthy, such systems must report well-calibrated confidence, meet explicit quality targets, and support appeals and transparency obligations that many jurisdictions now expect for high-stakes educational AI [euactannex, euactobl].
@@
Our design draws on prior evidence and seeks mathematics-specific insight. Building on our ASAG benchmark [kortemeyer2024performance] and reliability framework for handwritten mathematics [liu2024ai], we treat confidence calibration and routing as first-class objects: we integrate model-based signals (e.g., log probabilities, self-consistency), rubric alignment, and transcript/recognition quality into a calibrated probability of correctness, then set operating points that respect course and departmental constraints [kortemeyer2025assessing]. In parallel, we consider practical layout affordances linked to recognition and rubric alignment, echoing findings from physics and chemistry that small design choices can have outsized downstream effects [kortemeyer2024grading, cvengros2025assisting]. Throughout, we frame validity in terms of what matters for calculus learning: visibility into reasoning, representation changes (algebraic steps, graphs), and communication quality, not only terminal answers [meltzer2005relation, teodorescu2013new].
@@ 2.2 Sample
The student work was graded by teaching assistants (TAs) based on four to six rubric items per problem. For our study, we scanned 349 exams; Figure 1 shows a short excerpt as an example of graded student work. We were able to remove the red, pink, and green grading marks using the image-editing package OpenCV [opencv], as students were not allowed to use these colors, and we combined the pages for each problem, or, in case of problem 5, problem part; Figure 2 shows an example.
@@ 2.3 AI Grading
We accessed OpenAI models via Azure AI Services [azure] under the university’s privacy-preserving contractual framework. Specifically, we used the multimodal reasoning model GPT-5 [gpt5] (model version 2025-08-07) hosted in a Swedish data center. The code, including the prompt, is available at https://gitlab.ethz.ch/ethel/mathexam/-/blob/main/grade.py.
@@ 2.4 Confidence Filters
Generative AI will always produce an answer, regardless of whether that answer is reliable or not.. Thus, beyond computing AI-based scores, an equally important task is quantifying how much we should trust each score [kortemeyer2025assessing]. “Confidence” here differs from conventional quality metrics: in production use the ground truth is unknown, so confidence must be inferred from information available at decision time (the AI’s own outputs and model-based expectations). To do so, we draw on classical and Bayesian statistics.
@@ 2.4.1 Partial-credit Threshold
Evidence from earlier studies indicates that AI graders tend to be conservative — more prone to mark correct work as incorrect (false negatives) than the reverse (false positives) [kortemeyer2025assessing]. A simple, conservative safeguard is therefore to binarize partial credit using a threshold $t$ for “correctness.” For filtering purposes, any student-item with AI score below $t$ (i.e., incorrect or “insufficiently correct”) is flagged for human review, while only items at or above $t$ are considered for auto-acceptance (subject to the risk screen below).
@@ 2.4.2 Risk Threshold
where $\theta_{i}$ denotes the latent ability of student $i$, and $a_{j}$ and $b_{j}$ are the discrimination and difficulty of part $j$, respectively [rasch, kortemeyer2019quick].
Let $s_{ij}\in[0,1]$ be the AI’s normalized score for the same student-item. We define the risk of accepting that AI judgment as the absolute deviation between observed and expected [sinharay2006posterior]:
@@ 4 Discussion
The results are moderately promising but not perfect. Production-ready results can only be achieved for about 30% of the grading load, which is less than in our earlier studies of physics [kortemeyer2025assessing] and chemistry exams [cvengros2025assisting]. This might be improved by some practical, incremental adjustments that do not alter the mathematical substance of the exam:
@@ Acknowledgements
Ethical Considerations
This study was approved by the ETH Zurich Ethics Commission (IRB) as protocol EK 2023-N-366.
Consent to Participate
Consent to participate was not obtained for historical exam data, which was obtained within the framework of regular instruction (no experimental procedure). Per approved IRB protocol, data was depersonalized using a double-blind mechanism between original course instructors and staff carrying out the analysis.
Consent to Publish
Not applicable.